# GraphCheck: Breaking Long-Term Text Barriers with Extracted Knowledge Graph-Powered Fact-Checking

### v1

@@ 6 Conclusion
Training Data. Our method was trained on a 14k dataset with relatively short text. As shown in the paper, performance improves as the training data increases, indicating that there is still potential for further improvement if larger or higher-quality datasets are considered.
@@ A.3 Preprocessing for Benchmark
To extract graph information from the benchmark text data, we utilize LLM to separately extract entity-relation triples from both the claims and the documents. The extraction process follows the prompt shown in F. After preprocessing, the dataset is structured as {claim, doc, claim_kg, doc_kg, label}, where claim_kg and doc_kg represent the extracted KGs for the claim and document, respectively. Samples of the processed data are illustrated in Figure 8.
@@ Appendix B Synthetic Dataset for Training
To pre-train an external GNN, we synthesized a structured dataset of 14,000 samples based on the MiniCheck training set666We will release upon acceptance.. Using a method similar to A.3, we employed the Qwen2.5 7B model to extract KG triples from both the claim and document in each sample, following the prompt in F. Each sample is structured as {claim, doc, claim_kg, doc_kg, label}. Examples are shown in Figure 9.
@@ Appendix D Implementation Details
Appendix E Additional Experiments
@@
Appendix F Prompts
@@
Appendix G GraphRAG Evaluation
Implementation Details of GraphRAG. To streamline our implementation process, we leveraged the approach from the open-source nano-GraphRAG project999https://github.com/gusye1234/nano-graphrag for our testing phase. In our experiments, we followed the official approach by using GPT-4o and GPT-4o-mini models. GPT-4o was used for planning and generating responses, while GPT-4o-mini was used for summarization tasks. The workflow is divided into two phases: Insert and Query.
We compare GraphRAG and GraphCheck in Table 6. The results show that GraphRAG underperforms our proposed GraphCheck, which is based on Llama3.3 70B and Qwen 72B. Moreover, its cost is significantly higher than direct inference with GPT-4o (reaching $47.9 in Table 3’s scenario), yet the performance improvement is minimal. Given that GraphRAG is a general-purpose framework rather than one specifically designed for fact-checking, we suggest it may not be an ideal approach for this task.

### v2

@@ 6 Conclusion
Limited Training Data. The lack of high-quality training data is a common limitation in long-text fact-checking. For this reason, and to ensure a fair comparison, we adopted the 14K dataset provided by MiniCheck, which consists of relatively short texts. As shown in the paper, the performance of our method improves with increased training data, suggesting that there remains significant potential for further improvement if larger or higher-quality long-text datasets are considered.
Acknowledgments
Dr. Irene Li is supported by JST ACT-X (Grant JPMJAX24CU) and JSPS KAKENHI (Grant 24K20832). Dr. Qingyu Chen is supported by 1R01LM014604, National Library of Medicine, National Institutes of Health. This work used supercomputers provided by the Research Institute for Information Technology, Kyushu University, through the HPCI System Research Project (Project ID: hp250092). This work is also supported by NVIDIA Academic Grant Program and Google Cloud (Gemma 3 Academic Program).
@@ A.3 Preprocessing for Benchmark
To extract graph information from the benchmark text data, we utilize LLM to separately extract entity-relation triples from both the claims and the documents. The extraction process follows the prompt shown in H. After preprocessing, the dataset is structured as {claim, doc, claim_kg, doc_kg, label}, where claim_kg and doc_kg represent the extracted KGs for the claim and document, respectively. Samples of the processed data are illustrated in Figure 8.
@@ Appendix B Synthetic Dataset for Training
To pre-train an external GNN, we synthesized a structured dataset of 14,000 samples based on the MiniCheck training set. Using a method similar to A.3, we employed the Qwen2.5 7B model to extract KG triples from both the claim and document in each sample, following the prompt in H. Each sample is structured as {claim, doc, claim_kg, doc_kg, label}. Examples are shown in Figure 9.
@@ Appendix C Analysis of Computational Cost and Time Efficiency
Appendix D Additional Experiments
@@
Appendix E GraphRAG Evaluation
Implementation Details of GraphRAG. To streamline our implementation process, we leveraged the approach from the open-source nano-GraphRAG project666https://github.com/gusye1234/nano-graphrag for our testing phase. In our experiments, we followed the official approach by using GPT-4o and GPT-4o-mini models. GPT-4o was used for planning and generating responses, while GPT-4o-mini was used for summarization tasks. The workflow is divided into two phases: Insert and Query.
We compare GraphRAG and GraphCheck in Table 8. The results show that GraphRAG, based on GPT-4o, still underperforms our proposed GraphCheck, which is based on Llama3.3 70B and Qwen 72B. Fundamentally, GraphRAG is not the primary focus of our study. Although both involve graph structures, the core methods are different: GraphRAG relies on retrieval-based reasoning, while GraphCheck integrates structured knowledge directly into LLM for fact-checking. Additionally, GraphRAG is very costly, with expenses significantly exceeding those of direct inference with GPT-4o (reaching $47.9 in Table 3’s scenario), while offering only marginal performance improvements. Since GraphRAG is a general-purpose framework rather than one specifically designed for fact-checking, we believe it may not be the suitable approach for this task.
Appendix F Extended Discussion on How LLMs Understand Knowledge Graphs
Apart from fact-checking, using Knowledge Graphs (KGs) to enhance the text understanding ability of LLMs is also a popular research domain. A key challenge in this area is to enable LLMs to effectively understand graph-structured information. Existing methods Guo et al. (2023); Sen et al. (2023); Baek et al. (2023); Tang et al. (2024a) directly convert graph-structured information into natural language as part of the prompt, enabling LLMs to process graph knowledge without requiring additional graph encoders. However, these approach struggles to effectively handle complex KGs, often requiring multiple calls Sun et al. (2023) to help LLMs fully comprehend the structured information.
GNN-Based Methods. To address this limitation, existing methods Huang et al. (2024); Mavromatis and Karypis (2024) use Graph Neural Networks (GNNs) to encode complex KGs, allowing LLMs to focus on the most relevant entities and relationships. Specifically, G-Retriever He et al. (2024) uses a GNN to encode the retrieved KG as a graph embedding and converts the graph information into natural language as part of the prompt, feeding both into the LLM for better understanding. GNP Tian et al. (2024) encodes KGs into node embedding vectors and extracts the most relevant node embeddings, which are fed into the LLM along with text embeddings.
In this work, to enable efficient fact-checking on long-form texts, we use a pre-trained GNN to separately encode the entity relationships in the extracted claim KG and document KG. The resulting graph embeddings, combined with text prompt embeddings, enhance the LLM’s ability to understand complex logical structures in long texts. On the other hand, directly comparing the extracted document and claim KGs assists the LLM in performing fine-grained fact-checking within a single inference.
@@ Appendix F Extended Discussion on How LLMs Understand Knowledge Graphs
Appendix G Implementation Details
Appendix H Prompts